Write 3x^2 - 12x = 1 - y in quadratic function standard form,
y = mx^2 + bx + c

Answer by math_tutor2020(3835)


Standard form is y = ax^2 + bx + c

We need to move the 1 on the right hand side to the left side.
3x^2 - 12x = 1 - y
3x^2 - 12x - 1 = -y
-y = 3x^2 - 12x - 1

Then multiply everything by -1 to change the -y to y
-y = 3x^2 - 12x - 1
-1(-y) = -1(3x^2 - 12x - 1)
y = -3x^2 + 12x + 1 is the final answer.

The quadratic is now in standard form y = ax^2 + bx + c
a = -3
b = 12
c = 1

Standard form is useful if you want to apply the quadratic formula (as one example).